LadyRavyRay:

6 x 3 + 3^2 (The two is an exponent) Simplify

LadyRavyRay:

Help

snowflake0531:

Can you rewrite it-?

LadyRavyRay:

Um...I don't know....

supie:

I think she means \((6 \times 3) + 3^2\) ?

snowflake0531:

ohhh so what is 6 times 3 and what is 3 squared then add them together

LadyRavyRay:

Yes but without the pathesises

Shawnte:

they are needed

supie:

The parenthesis don't really matter lol 1st what is 3^2= ?

Shawnte:

in this equation we must use pemdas

Shawnte:

(6x3) + 3^2

LadyRavyRay:

Actually, in the equation I have, it's isn't shown.

Shawnte:

P E M D A S

Shawnte:

perithesis first

supie:

ladyravyray wrote:
Actually, in the equation I have, it's isn't shown.
the parenthesis ? or-

Shawnte:

then exponets

LadyRavyRay:

It says simplify the answer using the order of operations.

Shawnte:

then solve the addition problem to come

LadyRavyRay:

So I got 27?

supie:

I think that's right e.e

LadyRavyRay:

New question

LadyRavyRay:

@supie

LadyRavyRay:

@snowflake0531

LadyRavyRay:

@shawnte

supie:

1st exponents, 9^2= ?

LadyRavyRay:

18

Shawnte:

wsp

supie:

no, \(9^2=9\times9=?\)

Shawnte:

yep

LadyRavyRay:

oh yeah, 81

Shawnte:

yup

Shawnte:

now 81/3

LadyRavyRay:

27?

supie:

Yes, so now you have \(13+\frac{81}{3}\) 81/3= ?

Shawnte:

yup

Shawnte:

now 27+13

supie:

@ladyravyray wrote:
27?
no -

LadyRavyRay:

40?

Shawnte:

81/3.........yes it is

supie:

@ladyravyray wrote:
40?
yes e.e

Shawnte:

dh

supie:

ohh, i read that wrong

supie:

i thought she was saying 27 is the answer

LadyRavyRay:

It's ok.

Shawnte:

stop givin ppl wrong answeers den callin me wrong

supie:

@shawnte wrote:
stop givin ppl wrong answeers den callin me wrong
i didnt give anyone a wrong answer ?

supie:

when did i call u wrong

Shawnte:

@supie wrote:
@ladyravyray wrote:
27?
no -
thats acctually correct

supie:

27 isnt the answer

LadyRavyRay:

Thanks for the help y'all. I got the answer right with 40...

LadyRavyRay:

I got a 100% so, thanks!

supie:

np c:

Shawnte:

np and huh

supie:

shawnte wrote:
np and huh
"huh" what's up ? o:

LadyRavyRay:

Yeah, what is "huh" for?

supie:

dw abt it, some ppl are slow

LadyRavyRay:

True